OpenStudy (anonymous):

Harold has three dogs. The Irish terrier weighs one-third as much as the golden retriever. The golden retriever weighs half as much as the Saint Bernard. If Harold's Saint Bernard weighs one hundred fifty pounds, what is the combined weight of all three dogs? two hundred pounds two hundred twenty-five pounds two hundred fifty pounds two hundred seventy-five pounds

OpenStudy (anonymous):

GR = SB/2 GR = 150/2 GR = 75 IT = GR/3 IT = 75/3 IT = 25 SB + GR + IT = 150 + 75 + 25 = 250